OpenStudy (anonymous):

Find the value of y so that the line passing through the 2 points has a given slope. I would greatly appreciate it if you would show your work. (-1,5), (3,y) m=5 (5,7),(8,y), m=4/3 (2,-15),(5,y),m=4/5 Find the value of y so that the line passing through the 2 points has a given slope. I would greatly appreciate it if you would show your work. (-1,5), (3,y) m=5 (5,7),(8,y), m=4/3 (2,-15),(5,y),m=4/5 @algebra 1

jimthompson5910 (jim_thompson5910):

I'll do the first one to show you how to do them The slope formula is m = (y2-y1)/(x2-x1) Since the points are (-1,5), (3,y), we know that x1 = -1, y1 = 5, x2 = 3, and y2 = y So plug them all into the formula above to get m = (y - 5)/(3-(-1)) m = (y - 5)/(3+1) m = (y - 5)/4 Now plug in m = 5 (this is also given) to get 5 = (y-5)/4 Now solve for y 5 = (y-5)/4 5*4 = y-5 20 = y-5 20+5 = y 25 = y y = 25

OpenStudy (anonymous):

thx